News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 13 February 2004 9.30am EST

During the last week Chandra completed the observing schedule as planned.

A real-time procedure was completed on Feb 9 to update the on-board ephemeris.

A real-time procedure was run on Feb 12 to patch the OBC telemetry request list (TRL). The TRL was patched to allow the OBC to receive rate data from the HRC anti-coincidence and focal-plane detectors. This data is planned for future use within the radiation monitoring process, providing a backup to the EPHIN detector.

The schedule of targets for the next week is shown below and includes observations of IRAS 13349+2438 and NGC 1313 coordinated with HST.

------------------------------------------
IC2560 ACIS-S Feb 16 0313-192 ACIS-S Feb 17 Coma3 A ACIS-I Radiation Belts Feb 18 NGC 4217 ACIS-S Feb 19 GRO J2058+42 ACIS-I/HETG PKS 0208-512 ACIS-S Feb 20 MACS1108.8+0906 ACIS-I Radiation Belts Feb 21 FTM 0841+3604 ACIS-S NGC 1313 X-1 ACIS-S NGC 1313 X-2 ACIS-S Feb 22 IRAS 13349+2438 ACIS-S/HETG

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports